# Optimal Policy Minimum Bayesian Risk

### v1

@@ 5.4 Results Analysis
Figure 1 shows the comparison of the different generator-PRM pairs. The left shows pass@1 performance as a function of the real number of samples generated, averaged over all dataset examples. The right side shows study cases for specific optimal policy budgets of OPE-MBRD, signaled with a cyan star on the left side of the plot. Each marker on the right represents an example in the MATH-$500$ dataset, sorted from lower to higher difficulty. The difficulty is assessed by using the pass@1 of the normal generator $p(y\mid x)$ and the full $256$ samples. On the vertical axis we display the real number of samples $N$ used by OPE-MBRD averaged over experiment repetitions. We color as green instances for which the OPE-MBRD attains higher pass@1 than OP-MBRD, red if lower, black if both match (typically both reach $1.0$) and gray if both fail ($0.0$). We consider generator-PRM pair as calibrated if the number of samples used increases with problem difficulty and this leads to performance improvements (green dot).
@@ Appendix B Development Set Results
Figure 3: Development set (Numinamath-$500$) results.
As stated in Section 5.3, we created a dev set for hyperparameter tuning based on NuminaMath [^13] https://huggingface.co/datasets/AI-MO/NuminaMath-CoT of the same size as MATH-$500$. We report full results on the development set under the same conditions as the MATH-$500$ test set in Figure 3. These results were used to tune $beta$ for different generator/PRM pairs. As it can be observed from the results, this dataset is harder than Math-$500$, but model/PRM calibration is similar. Overall improvements with OP-MBRD are also larger, but this can be due to tuning effects.
Appendix C Hardware
@@
Appendix D Full Derivation
@@ D.1 Optimal Policy
due to the properties of the KL divergence, $\mathcal{L}(q)$ is maximized ($=0$) when the two distributions in the $\mathrm{KL}$ match i.e.
@@
D.2 Rao-Blackwellized Rejection Sampling
@@
D.3 Importance Sampling Equivalence and Expected Number of Optimal Policy Samples

### v2

@@ 5.4 Results Analysis
Figure 1 shows the comparison of the different generator-PRM pairs. The left shows pass@1 performance as a function of the real number of samples generated, averaged over all dataset examples. The right side shows study cases for specific optimal policy budgets of OPE-MBRD, signaled with a cyan star on the left side of the plot. Each marker on the right represents an example in the MATH-$500$ dataset, sorted from lower to higher difficulty. The difficulty is assessed by using the pass@1 of the normal generator $p(y\mid x)$ and the full $256$ samples. On the vertical axis we display the real number of samples $N$ used by OPE-MBRD averaged over experiment repetitions. We color as green instances for which the OPE-MBRD attains higher pass@1 than OP-MBRD, red if lower, black if both match (typically both reach $1.0$) and gray if both fail ($0.0$). We consider generator-PRM pair as well calibrated if the number of samples used increases with problem difficulty and this leads to performance improvements (green dot). Improvements below $0.02$ are ignored to avoid noise.
@@ Appendix B Development Set Results
Figure 3: Development set (Numinamath-$500$ subset) results. Left: pass@1 score as a function of the number of samples per input with average and standard deviation displayed. A dashed line denotes the efficient version (OPE-MBRD). A star marker denotes the OPE-MBRD configuration represented on the right side. Right: Number of samples OPE-MBRD selects for every example in the test set, sorted from easy to difficult by regular decoding pass@1.
As stated in Section 5.3, we created a dev set for hyperparameter tuning based on NuminaMath [^14] https://huggingface.co/datasets/AI-MO/NuminaMath-CoT of the same size as MATH-$500$. We report full results on the development set under the same conditions as the MATH-$500$ test set in Figure 3. These results were used to tune betabeta for different generator/PRM pairs. As it can be observed from the results, this dataset is harder than Math-$500$, but model/PRM calibration is similar. Overall improvements with OP-MBRD are also larger, but this can be due to tuning effects.
Appendix C Effect of Reference Policy
Figure 4: Effect of the reference policy on OP-MBRD. Experimental setup is that of Figure 1 for Qwen-2.5-1.5b-math-Instruct using Qwen-2.5-72b-math-Instruct as reference policy. Note that OP-MBRD (+log-ratio) overlaps with OP-MBRD
Figure 4 shows the effect of the log-ratio component on the overall OP-MBRD. In Equation 10, this amounts to setting the reference policy equal to the generator (OP-MBRD), equal to a given reference policy (OP-MBRD + log-ratio) or the latter but setting the reward to $1.0$ (log-ratio only). As explained in Section 4.4 the last one should be equivalent to self-normalized importance sampling from the reference model. As reference policy, we use Qwen-2.5-72b-math-Instruct. Different values of $\beta$ were tried in the dev set. Log-ratio alone showed no sensitivity to $\beta$ while OP-MBRD + log-ratio resulted in same or degraded performance. We present here results on the math-$500$ test set. The nominal values for the experimental setup of Section 5 are used here. The log-ratio component alone does show improvements over conventional MBRD without the need for a PRM, but these decrease with model size. When combined with a PRM, the improvement overall is negligible.
Appendix D Hardware
@@
Appendix E Full Derivation
@@ E.1 Optimal Policy
due to the properties of the KL divergence, $\mathcal{L}(q)$ is maximized (=0=0) when the two distributions in the $\mathrm{KL}$ match i.e.
@@
E.2 Rao-Blackwellized Rejection Sampling
@@
E.3 Importance Sampling Equivalence and Expected Number of Optimal Policy Samples